OpenStudy (anonymous):

marina drove 288 miles in the same amount of time it took Jaime drive 312 miles if Jaime drove 6 miles per hour faster than marina at which speeds were Jaime and marina traveling?

OpenStudy (johnweldon1993):

Hmm Marina drove some rate for some time and went 288 miles Jaime drove at the same rate plus 6mph for the same amount of time and went 312 miles \[\large Marina \rightarrow t = \frac{D}{r} \] \[\large Jaime \rightarrow t = \frac{D}{r}\] We know their times were equal...so we can set those 2 equations equal to each other \[\large \frac{D_m}{r_m} = \frac{D_j}{r_j}\] Now we just plug in what we know \[\large \frac{288}{x} = \frac{312}{x + 6}\] And we cross multiply and simplify

OpenStudy (johnweldon1993):

\[\large 288(x + 6) = 312x\] \[\large 288x + 1728 = 312x\] \[\large 24x = 1728\] \[\large x = 72\] So that means \[\large Marina = 72mph\] And \[\large Jaime = 72 + 6 = 78mph\]

OpenStudy (anonymous):

thanks I need one more help

OpenStudy (anonymous):

mason lives in highland and works in ridgeview there are 5 roads from highland to ridgrview and one of the roads is a one-way street he wants to find the number of different round trips he can take to and from work he plans to drives on each street no more than 3time during all of his round trips . how many different round trips can mason take to and from work